### v2

@@ 2.1.3 Circular Dichroism and UV Absorption Spectroscopy
Circular dichroism (CD) is a widely used technique in biochemistry for studying the secondary structure of biomolecules. By examining the ellipticity $\mathrm{\theta\lambda}$, which is proportional to the differential absorption of circularly polarized light, one can characterize how the chiral structures of proteins, such as alpha helices and beta barrels, change under various conditions. At the same time, the UV absorbance spectra serve as a complementary tool providing important information about protein stability, unfolding, and aggregation.
@@
We observe that at approximately 52C in both the CD/UV absorbance spectra of free Fib and Fib with SiO2NPs, there is a noticeable modification of the experimental profiles. Specifically, they indicate i) a decrease in UV absorption (Fig. 6 a and e), and ii) a change in Fib’s secondary structure (Fig. 6 b and f of the CD spectra). These changes correspond to an increase in the intensity of the water band in the Raman spectra for these two cases (Fig. 4 a and c).
On the other hand, the UV absorption of Fib with CNPs shows no $T$-dependence (Fig. 6 c), while the CD spectra, which are noisy due to CNP’s absorption, exhibit a $T$-dependence that is weaker than those observed for free Fib and Fib with SiO2NPs (Fig. 6 d). These data are only partially consistent with the corresponding results from UVRR spectra, displaying a very weak dependence of the spectra on temperature (Fig. 4 b).
@@
Therefore, as with the UVRR spectra, while we observe differences and similarities among the three cases, no quantitative results can be systematically extracted from these data. To achieve an analysis of the spectroscopic data that quantitatively evaluates how far the NP-adsorbed proteins are from their native state, we have developed a method based on ML approach for the combined spectra.
Figure 6: Temperature-dependent UV absorption and CD spectra collected for Fib 2 mg/ml under various conditions. In all panels, the temperature ranges between 22 and 88C. Upper panels a), c), and e) are the UV absorption spectra for Fib in PBS pH 7.4, Fib with CNPs, and Fib with SiO2NPs, respectively. Lower panels b), d), and f) are the corresponding CD spectra. Curves with different colors correspond to different temperatures as indicated by the upper color bar. Dashed lines in panels c)-f) represent the spectra of panels a) and b) at the lowest temperature, provided here as a reference. Panels g) and h) display the UV absorbance at 225 nm, and the ellipticity $\mathrm{\theta\lambda}$, measured in millidegrees (mdeg), at 222 nm, respectively, for the three cases: free Fib in PBS (black circles), Fib with CNPs (red pentagons), and Fib with SiO2NPs (green squares).
@@ 2.2 Machine Learning Automate Workflow for High-dimensional Composite Spectra
Finally, we apply the ML model, which includes (i) metric analysis, (ii) clustering, manifold reduction, and (iii) the subsequent elaboration of the outputs. Different tools, such as Euclidean and Manhattan distances, along with PCA, are routinely used to analyze high-dimensional and spectral data. For example, the distances can measure the similarity between pairs of spectra to quantify the structural changes of proteins, while PCA can visualize the patterns of the original data spectra in a two-dimensional space. However, these traditional tools encounter issues due to the variety and high dimensionality of the data at hand. For this reason, we replace them with the Wasserstein distance and t-SNE, respectively, as discussed in the next section. Our entire work pipeline, starting from the experiments and concluding with the ML analysis, is illustrated in Fig. 7.
@@ 3 ML model
Therefore, the impact of the NP surface is substantial, inducing only a gradual structural change. This analysis enables us to better interpret the original spectra, clarifying that the structural disorder induced by the increase in temperature in this case does not resemble the standard unfolding that occurs for the free protein.
These observations regarding the differences in the Wasserstein distances of Fib proteins interacting with hydrophilic silica NPs and hydrophobic carbon NPs indicate that the two interfaces impact Fib’s structure in fundamentally different ways. Furthermore, recalling that we normalize the Wasserstein distance relative to the minimum and maximum disorder in the free Fib case, Fig. 8 provides information about the absolute value of disorder as a function of $T$ in each case. In particular, Fib on silica NPs unfolds approximately 90% of the free case, while Fib on carbon NPs is about 50% folded at the same $T$.
It is worth noting that using Euclidean and Manhattan distances would result in a different erroneous scenario, where several substantial changes, even at the lowest temperatures, appear across the temperature range. This is discussed in Section 2 (Figs. S2a, S2b) of ESI.
Figure 8: Normalized Wasserstein distance computed for composite spectra of free Fib and Fib with NPs. Upper panel: free Fib (squares). Central panel: Fib with CNPs (circles). Lowe panel: Fib with SiO2NP (triangles). In each case, the quantity is normalized with respect to the values at $22^{\circ}$ and $86^{\circ}$ for the free case, corresponding to Fib in its native (folded) or disordered (unfolded) state. Therefore, it quantifies the relative distance of the spectra (structural disorder) from the reference spectrum of a free Fib protein in its native state at $22^{\circ}$.
@@
To compute t-SNE, an input parameter called the perplexity $\mathcal{P}$ is required. The perplexity roughly corresponds to the effective number of neighbors in the high-dimensional space for each point considered when calculating similarities. In cases where several thousand data sets are available, the suggested default values for $\mathcal{P}$ are between 30 and 50 [30]. However, in our case, due to the small size of each dataset with only 17 temperatures, we propose to automatically find the optimal value of $\mathcal{P}$ through a grid search based on K-means clustering, a popular unsupervised ML algorithm used for clustering data, and the average silhouette score, which gives an overall measure of how appropriately the data has been clustered, as discussed in Section 4 of ESI [58, 60]. According to our procedure, the perplexity is set to 7, consistent with the small size of the datasets. Other parameters are set to their default values for t-SNE optimization [30].
@@
Figure 9: t-SNE analysis of composite spectra. a) Two-dimensional embedding from t-SNE with cosine similarity of composite spectra of free Fib (squares), Fib with CNPs (circles), and Fib with SiO2NPs (triangles), each corresponding to a given temperature. Temperatures are indicated in the color bar, with $\Delta T=4^{\circ}$ starting from $T=22^{\circ}$. We used $\mathcal{P}=7$ and PCA initialization for t-SNE computations. b) Same as in panel a), with centroids (red cross marks) computed using the K-means clustering algorithm, assuming the optimal number of clusters $n_{\rm op}=5$ as input. In this case, the average silhouette score is $\approx 0.94$.
@@
Although not quantitative, the t-SNE analysis qualitatively confirms that the structural change for Fib with SiO2NPs is less significant than in the free Fib case, even though the Fib unfolds at the same temperature in both scenarios. In contrast, it highlights the gradual increase of disorder without unfolding for Fib adsorbed on the hydrophobic carbon NPs up to $86^{\circ}$.
We emphasize that if PCA is used instead of t-SNE, the resulting two-dimensional projection would suggest the presence of five or more clusters. Furthermore, the calculation of the centroids would be less clear, making it more difficult to quantitatively evaluate the absolute amount of structural disorder in the three cases. This is because PCA identifies some data points as outliers, as discussed in Section 3 of the ESI, due to its difficulty in handling data originating from very different sources.
@@ 4 Conlcusion
We developed a method that leverages ML to interpret spectral data from multiple experimental techniques, measuring the structural properties of proteins. This approach optimizes the combined information and systematically evaluates the relative and absolute structural disorder of proteins as a function of varying external conditions.
@@
As a relevant case study, we consider the debated effect of NP surface chemistry on the structural changes of proteins adsorbed within the corona at varying temperatures. To achieve this goal, we consider fibrinogen at blood concentrations when interacting with hydrophobic carbon or hydrophilic silicon dioxide nanoparticles. Our findings demonstrate that the NP chemistry has a significant impact on the stability of Fib as a function of temperature, with striking differences between the hydrophilic and hydrophobic cases.
@@ Acknowledgments
All authors acknowledge Elettra Sincrotrone Trieste for providing access to its synchrotron radiation facilities and for financial support under the IUS internal project. A.M.-S. and M.P.M. gratefully acknowledge the financial support through the H2020 SUNSHINE grant no. 952924. G.F. and G.M. acknowledge the support by MCIN/AEI/ 10.13039/ 501100011033, the European Commission “ERDF A way of making Europe” grant number PID2021-124297NB-C31, and the support by ICREA Foundation (ICREA Academia prize). G.F. also acknowledges the support from the Ministry of Universities 2023-2024 Mobility Subprogram within the Talent and its Employability Promotion State Program (PEICTI 2021-2023), the Visitor Program of the Max Planck Institute for the Physics of Complex Systems for supporting a visit started in November 2022, the COST Action CA22143 EuMINe. G.M. would like to thank Jozsef Kardos for his help in using the BeStSel software. All authors thank Oriol Miró for testing the code.
@@
A.M-S. performed the experimental measurements and supported the design and implementation of the analysis. G.M. designed and implemented the Machine Learning pipeline/protocol. F.D. and B.R. provided support for the spectroscopic measurements. B.R. coordinated the activity at Elettra Sincrotrone. I.F. provided the nanoparticles. A.M.-S. and G.M. wrote the first version of the manuscript. A.M.-S., G.M., and B.R. processed the data and prepared the figures. A.M-S., G.M., B.R., M.M., and G.F. discussed the analysis, the strategies for their elaboration, and the interpretation of the results. G.M., A.M-S. and G.F. discussed the design of the computational method and the strategies for its implementation. G.M. prepared the support webpage repository for the open-source codes. B.R., M.M., and G.F. prepared revised versions of the manuscript. A.M-S. wrote the final draft. All the authors discussed and contributed to the final version of the manuscript. M.M. and G.F. acquired funds, designed the study, and coordinated the experimental and theoretical activities.
@@ Appendix A Materials
SiO2NPs of 0.1µm (stock concentration of 50 mg/mL) were provided by Kisker Biotech GmbH (Germany). CNPs were synthesized in the lab. Phosphate buffer saline (PBS) tablets, Eppendorf LoBind microcentrifuge tubes, Fibrinogen from human plasma (F3879) and D-(+)-Sucrose (99.9%) were purchased from Sigma-Aldrich (Ireland). One PBS tablet was dissolved in 200 mL of ultrapure water to obtain a 0.01 M phosphate buffer, 0.0027 M potassium chloride and 0.137 M sodium chloride solution (pH 7.4 at 25C). TEM grids of Formvar / Silicon monoxide 200 mesh with copper approximate grid hole size 97µm were purchased from Ted Pella (USA).
@@ B.1.1 Carbon nanoparticle synthesis
Carbon nanoparticles of 100 nm nominal diameter were synthesised using the method developed by Kokalari et al. [61] They were produced by carbonization of glucose in water, consisting mainly of amorphous carbon with graphitic residues incorporated in the amorphous matrix. Compared to other carbon nanomaterials, they have the advantage of being produced by a synthesis that does not involve toxic solvents, with a high yield and, moreover, with a great control of the dimensions having a low variability from lot to lot. They also have a functionalized surface that makes them stable and negatively charged in water thanks to the acidic carboxylic groups exposed on the surface.
@@ B.1.2 Nanoparticle characterization
Dynamic Light Scattering (DLS) measurements were performed using Zetasizer Nano ZS (Malvern). The sample cuvettes were equilibrated at 25C for 90 s. For each measurement, the number of run and duration were automatically determined and repeated three times. Data analysis has been performed according to standard procedures and interpreted through a cumulant expansion of the field autocorrelation function to the second order.
@@
Transmission electron microscopy (TEM) images were obtained with a JEOL JEM-1400PLUS transmission electron microscope operating at an acceleration voltage of 120 kV. The statistical analysis was carried out using ImageJ and average size and standard deviation was measured on 103 NPs.
@@ B.1.3 Protein Corona obtention
[…]
@@ B.2 UV Resonant Raman measurements
UVRR spectra were collected by using the synchrotron-based UVRR set-up available at the BL10.2-IUVS beamline of Elettra Sincrotrone Trieste (Italy). [62] The excitation wavelength at 226 nm was obtained by setting the energy of the emission of the Synchrotron Radiation (SR) light by regulating the undulator gap aperture. The incoming SR light was further monochromatized through a Czerny-Turner monochromator (750 nm focal length Acton SP2750, Princeton Instruments, Acton, MA, USA) equipped with a holographic grating with 3600 grooves/mm. The Raman signal of the protein solutions was collected in back-scattered geometry and analyzed via a single-pass Czerny-Turner spectrometer (Trivista 557, Princeton Instruments, 750 mm of focal length) equipped with a holographic grating at 1800 g/mm. A thermoelectric cooled CCD camera optimized for the UV range was used to detect the signal at the output of the spectrometer. The resolution was set at about 1.6 $\mathrm{cm^{-1}}$/pixel. Cyclohexane (spectroscopic grade, Sigma Aldrich) was used for the calibration of the wavenumber scale. The final radiation power on the samples was about 20 $\mathrm{\mu}$W. To avoid photodegradation during the measurements, the samples were continuously oscillated with 1 Hz frequency and a path length of 1 mm. For the collection of the temperature-dependent measurements (in the temperature range 295-359 K with step of 4 K) on protein solutions a sample holder equipped with a thermal bath coupled to a resistive heating system was used to control the temperature with a stability of ± 0.1 K. The solution of protein with and without NP were measured at concentration of 2 mg/ml (the same weight/molar concentration was kept for protein and NP).

### v3

@@ 2.1.3 Circular Dichroism and UV Absorption Spectroscopy
Circular dichroism (CD) is a widely used technique in biochemistry for studying the secondary structure of biomolecules. By examining the ellipticity $\mathrm{\theta_{\lambda}}$, which is proportional to the differential absorption of circularly polarized light, one can characterize how the chiral structures of proteins, such as alpha helices and beta barrels, change under various conditions. At the same time, the UV absorbance spectra serve as a complementary tool providing important information about protein stability, unfolding, and aggregation.
@@
We observe that at approximately 52C in both the CD/UV absorbance spectra of free Fib and Fib with SiO2NPs, there is a noticeable modification of the experimental profiles. Specifically, they indicate i) a decrease in UV absorption (Fig. 6 a and e), and ii) a change in Fib’s secondary structure (Fig. 6 b and f of the CD spectra). In particular, the flattening of the CD spectra indicates a loss of ellipticity, which is sensitive to the protein’s secondary structure, and the decrease in intensity at 208 nm and 222 nm suggests a loss of $\alpha$-helical moieties. These changes correspond to an increase in the intensity of the water band in the Raman spectra for these two cases (Fig. 4 a and c).
On the other hand, the UV absorption of Fib with CNPs shows no $T$-dependence (Fig. 6 c), while the CD spectra, which are noisy due to CNP’s absorption, exhibit a $T$-dependence that is weaker than those observed for free Fib and Fib with SiO2NPs (Fig. 6 d). These data are consistent with the corresponding results from UVRR spectra, displaying a very weak dependence of the spectra on temperature (Fig. 4 b).
@@
Therefore, as with the UVRR spectra, while we observe differences and similarities among the three cases, no quantitative results can be systematically extracted from these data, as further discussed in Section 3 (Fig. S3) of the ESI. To achieve an analysis of the spectroscopic data that quantitatively evaluates how far the NP-adsorbed proteins are from their native state, we have developed a method based on an ML approach for the combined spectra.
Figure 6: Temperature-dependent UV absorption and CD spectra collected for Fib 0.25 mg/mL under various conditions. In all panels, the temperature ranges between 22 and 88C. Upper panels a), c), and e) are the UV absorption spectra for Fib in PBS pH 7.4, Fib with CNPs, and Fib with SiO2NPs, respectively. Lower panels b), d), and f) are the corresponding CD spectra. Curves with different colors correspond to different temperatures as indicated by the upper color bar. Dashed lines in panels c)-f) represent the spectra of panels a) and b) at the lowest temperature, provided here as a reference. Panels g) and h) display the UV absorbance at 225 nm, and the ellipticity $\mathrm{\theta\lambda}$, measured in millidegrees (mdeg), at 222 nm, respectively, for the three cases: free Fib in PBS (black circles), Fib with CNPs (red pentagons), and Fib with SiO2NPs (green squares).
@@ 2.2 Machine Learning Automate Workflow for High-dimensional Composite Spectra
Finally, we apply the ML model, which includes (i) metric analysis, (ii) clustering, manifold reduction, and (iii) the subsequent elaboration of the outputs. Different tools, such as Euclidean and Manhattan distances, along with PCA, are routinely used to analyze high-dimensional and spectral data. For example, the distances can measure the similarity between pairs of spectra to quantify the structural changes of proteins. At the same time, PCA can visualize the patterns of the original data spectra in a two-dimensional space. However, these traditional tools encounter issues due to the variety and high dimensionality of the data at hand. For this reason, we replace them with the Wasserstein distance and t-SNE, respectively, as discussed in the next section. Our entire work pipeline, starting from the experiments and concluding with the ML analysis, is illustrated in Fig. 7.
@@ 3 ML model
Therefore, the impact of the NP surface is substantial, inducing only a gradual structural change. This analysis enables us to interpret the original spectra better, clarifying that the structural disorder caused by the increase in temperature in this case does not resemble the standard unfolding that occurs for the free protein.
These observations regarding the differences in the Wasserstein distances of Fib proteins interacting with silica NPs and carbon NPs indicate that the two interfaces impact Fib’s structure in fundamentally different ways. Furthermore, recalling that we normalize the Wasserstein distance relative to the minimum and maximum disorder in the free Fib case, Fig. 8 provides information about the absolute value of disorder as a function of $T$ in each case. In particular, Fib on silica NPs unfolds approximately 90% of the free case, while Fib on carbon NPs is about 50% folded at the same $T$.
It is worth noting that using Euclidean and Manhattan distances would result in a different erroneous scenario, where several substantial changes, even at the lowest temperatures, appear across the temperature range. This is discussed in Section 1 of ESI.
Figure 8: Normalized Wasserstein distance computed for composite spectra of free Fib and Fib with NPs. Upper panel: free Fib (squares). Central panel: Fib with CNPs (circles). Lower panel: Fib with SiO2NP (triangles). In each case, the quantity is normalized with respect to the values at $22^{\circ}$ and $86^{\circ}$ for the free case, corresponding to Fib in its native (folded) or disordered (unfolded) state. Therefore, it quantifies the relative distance of the spectra (structural disorder) from the reference spectrum of a free Fib protein in its native state at $22^{\circ}$.
@@
To compute t-SNE, an input parameter called the perplexity $\mathcal{P}$ is required. As discussed in Section 4 of ESI, the perplexity is set to 7, consistent with the small size of the datasets. Other parameters are set to their default values for t-SNE optimization [33].
@@
Figure 9: t-SNE analysis of composite spectra. a) Two-dimensional embedding from t-SNE with cosine similarity of composite spectra of free Fib (squares), Fib with CNPs (circles), and Fib with SiO2NPs (triangles), each corresponding to a given temperature between $22^{\circ}$ and $86^{\circ}$, at $4^{\circ}$ intervals. The temperature of each data point is color-coded as indicated in the vertical color bar. . b) Same as in panel a), with centroids (red cross marks) computed using the K-means clustering algorithm, with average silhouette score $\approx 0.94$. In both panels, we use $\mathcal{P}=7$, optimal number of clusters $n_{\rm op}=5$, and PCA initialization as discussed in Section 4 of ESI.
@@
Although not quantitative, the t-SNE analysis qualitatively confirms that the structural change for Fib with SiO2NPs is less significant than in the free Fib case, even though the Fib unfolds at the same temperature in both scenarios. In contrast, it highlights the gradual increase of disorder without unfolding for Fib adsorbed on the carbon NPs up to $86^{\circ}$.
We emphasize that if PCA is used instead of t-SNE, the resulting two-dimensional projection would suggest the presence of five or more clusters. Furthermore, the calculation of the centroids would be less precise, making it more difficult to quantitatively evaluate the absolute amount of structural disorder in the three cases. This is because PCA identifies some data points as outliers, as discussed in Section 2 of the ESI, due to its difficulty in handling data originating from very different sources.
@@ 4 Conlcusion
We developed a method that leverages ML to interpret spectral data from multiple experimental techniques, measuring proteins’ structural properties. This approach optimizes the combined information and systematically evaluates proteins’ relative and absolute structural disorder as a function of varying external conditions.
@@
As a relevant case study, we consider the debated effect of NP surface chemistry on the structural changes of proteins adsorbed within the corona at varying temperatures. To achieve this goal, we consider fibrinogen at blood concentrations when interacting with partially hydrophobic carbon or hydrophilic silicon dioxide nanoparticles. Our findings demonstrate that the NP chemistry has a significant impact on the stability of Fib as a function of temperature, with striking differences between the two cases.
@@ Acknowledgments
All authors acknowledge Elettra Sincrotrone Trieste for providing access to its synchrotron radiation facilities and for financial support under the IUS internal project. A.M.-S. and M.P.M. gratefully acknowledge the financial support through the H2020 SUNSHINE grant no. 952924. G.F. and G.M. acknowledge the support by Grant PID2021-124297NB-C31 funded by Spanish MICIU/AEI/10.13039/501100011033 and European Union “ERDF A way of making Europe”. G.F. also acknowledges the support from the Ministry of Universities 2023-2024 Mobility Subprogram within the Talent and its Employability Promotion State Program (PEICTI 2021-2023), the Visitor Program of the Max Planck Institute for the Physics of Complex Systems for supporting a visit started in November 2022, the COST Action CA22143 EuMINe. G.M. would like to thank Jozsef Kardos for his help in using the BeStSel software. All authors thank Oriol Miró for testing the code.
@@
A.M.-S. performed the experimental measurements and supported the design and implementation of the analysis. G.M. designed and implemented the Machine Learning pipeline/protocol. F.D. and B.R. provided support for the spectroscopic measurements. B.R. coordinated the activity at Elettra Sincrotrone. I.F. provided the nanoparticles. A.M.-S. and G.M. wrote the first version of the manuscript. A.M.-S., G.M., and B.R. processed the data and prepared the figures. A.M.-S., G.M., B.R., M.M., and G.F. discussed the analysis, the strategies for their elaboration, and the interpretation of the results. G.M., A.M-S., and G.F. designed the computational method and the plan for its implementation. G.M. prepared the support webpage repository for the open-source codes. B.R., M.M., and G.F. prepared revised versions of the manuscript. A.M.-S. wrote the final draft. All the authors discussed and contributed to the final version of the manuscript. M.M. and G.F. acquired funds, designed the study, and coordinated the experimental and theoretical activities.
@@ Appendix A Materials
SiO2NPs of 0.1µm (stock concentration of 50 mg/mL) were provided by Kisker Biotech GmbH (Germany). CNPs were synthesized in the lab. Phosphate buffer saline (PBS) tablets, Eppendorf LoBind microcentrifuge tubes, Fibrinogen from human plasma (F3879) and D-(+)-Sucrose (99.9%) were purchased from Sigma-Aldrich (Ireland). One PBS tablet was dissolved in 200 mL of ultrapure water to obtain a 0.01 M phosphate buffer, 0.0027 M potassium chloride, and 0.137 M sodium chloride solution (pH 7.4 at 25C). TEM grids of Formvar / Silicon monoxide 200 mesh with copper approximate grid hole size of 97 $\mathrm{\mu}$m, were purchased from Ted Pella (USA).
@@ B.1.1 Carbon nanoparticle synthesis
Carbon nanoparticles of 100 nm nominal diameter were synthesised using the method developed by Kokalari et al. [66]. They were produced by the carbonization of glucose in water, consisting mainly of amorphous carbon with graphitic residues incorporated in the amorphous matrix. Compared to other carbon nanomaterials, they have the advantage of being produced with precise control of the dimensions [67]. Moreover, they exhibit negatively charges on the surface at neutral pH, due to the presence of acidic carboxylic or phenolic groups exposed on the surface, which allow one to tune the hydrophobicity of the NPs.
@@ B.1.2 Nanoparticle characterization
Dynamic Light Scattering (DLS) measurements were performed using Zetasizer Nano ZS (Malvern). The sample cuvettes were equilibrated at 25C for 90 s. For each measurement, the number of runs and duration were automatically determined and repeated three times. Data analysis was performed according to standard procedures and interpreted through a cumulant expansion of the field autocorrelation function to the second order (Fig. S6 in ESI).
@@
Transmission electron microscopy (TEM) images were obtained with a JEOL JEM-1400PLUS transmission electron microscope operating at an acceleration voltage of 120 kV. The statistical analysis was carried out using ImageJ, and the average size and standard deviation were measured on 103 NPs.
@@ B.1.3 Protein Corona obtention
CNPs from a 4 mg/mL stock solution and SiO2NPs from a 50 mg/mL stock solution were introduced into low protein-binding 1.5 mL microtubes and diluted to 2 mg/mL, respectively. We weighed 4 mg of Fib using a microbalance and diluted it in PBS to prepare the protein stock solution, which was later transferred into the NP solution to achieve a diluted concentration of 2 mg/mL. Consequently, the NP-protein corona complexes were formed.
A total volume of 500 $\mathrm{\mu}$L from the final solution was transferred to a 1 mm cell to perform the UVRR. For the DCS, 100 $\mathrm{\mu}$L from the final sample were used at each run. For the DLS (Fig. S7 in ESI), the in-situ corona sample was diluted ten times to get an optimal attenuation. For the CD and UV absorption, the sample was diluted four times to avoid saturation of the signal.
@@ B.2 UV Resonant Raman measurements
UVRR spectra were collected by using the synchrotron-based UVRR set-up available at the BL10.2-IUVS beamline of Elettra Sincrotrone Trieste (Italy). [68] The excitation wavelength at 226 nm was obtained by setting the energy of the emission of the Synchrotron Radiation (SR) light by regulating the undulator gap aperture. The incoming SR light was further monochromatized through a Czerny-Turner monochromator (750 nm focal length Acton SP2750, Princeton Instruments, Acton, MA, USA) equipped with a holographic grating with 3600 grooves/mm. The Raman signal of the protein solutions was collected in back-scattered geometry and analyzed via a single-pass Czerny-Turner spectrometer (Trivista 557, Princeton Instruments, 750 mm of focal length) equipped with a holographic grating at 1800 g/mm. A thermoelectric cooled CCD camera optimized for the UV range was used to detect the signal at the output of the spectrometer. The resolution was set at about 1.6 $\mathrm{cm^{-1}}$/pixel. Cyclohexane (spectroscopic grade, Sigma Aldrich) was used for the calibration of the wavenumber scale. The final radiation power on the samples was about 20 $\mathrm{\mu}$W. To avoid photodegradation during the measurements, the samples were continuously oscillated with a 1 Hz frequency and a path length of 1 mm. For the collection of the temperature-dependent measurements (in the temperature range 22-86C with a step of 4 C) on protein solutions, a sample holder equipped with a thermal bath coupled to a resistive heating system was used to control the temperature with a stability of ±$\pm 0.1^{\circ}$. The solution of protein with and without NP was measured at a concentration of 2 mg/mL (the same weight/molar concentration was kept for protein and NP).